tan(A) is the integral solution of inequality 4x^2-16x+15

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

4x^2 - 16x + 15 = 0
4x^2 - 10x - 6x + 15 = 0
2x(2x - 5) - 3(2x - 5) = 0
(2x - 3)(2x - 5) = 0
x = \frac{3}{2}, \frac{5}{2}
tanA = \frac{3}{2}, \frac{5}{2}
A = tan^{-1}\frac{3}{2}, tan^{-1}\frac{5}{2}